Find the velocity of a transverse wave in a string which is 4.0 m long with a


total mass of 20 g when under a tension of 100N.


1
Expert's answer
2021-12-20T10:31:07-0500

We can find the velocity of a transverse wave in a string from the formula:


"v=\sqrt{\dfrac{T}{\mu}}=\sqrt{\dfrac{TL}{m}},""v=\sqrt{\dfrac{100\ N\times4\ m}{0.02\ kg}}=141.4\ \dfrac{m}{s}."